Express the roots of the equation z3 − α3 = 0 in terms of α and w, where w is a complex cube root of unity. Use your answer to find the roots of the following equations in the form a + ib.
1
Expert's answer
2020-05-20T19:38:34-0400

Solution.

"z^3-\alpha^3=0;"

"z^3-\alpha^3=(z-\alpha)(z^2+\alpha z+\alpha^2);"

"(z-\alpha)(z^2+\alpha z+\alpha^2)=0;"

"z-\alpha=0" or "z^2+\alpha z+\alpha^2=0;"

"z_1=\alpha;" "D=\alpha^2-4\alpha^2=3\alpha^2;"

"z_2=\dfrac{-\alpha+\sqrt{3\alpha^2}}{2}=\alpha \dfrac{-1+i\sqrt{3}}{2};"

"z_3=\dfrac{-\alpha-\sqrt{3\alpha^2}}{2}=\alpha \dfrac{-1-i\sqrt{3}}{2};"

Cube Root of Unity Value:

"\omega_1=1" - real;

"\omega_2= \dfrac{-1+i\sqrt{3}}{2}" - complex;

"\omega_3= \dfrac{-1-i\sqrt{3}}{2}" - complex;

"z_1=\alpha\sdot\omega_1=\alpha\sdot1=\alpha;"

"z_2=\alpha\sdot\omega_2=\alpha \dfrac{-1+i\sqrt{3}}{2}=-\dfrac{1}{2}\alpha+i\dfrac{\alpha\sqrt{3}}{2};"

"z_3=\alpha\sdot\omega_3=\alpha \dfrac{-1-i\sqrt{3}}{2}=-\dfrac{1}{2}\alpha-i\dfrac{\alpha\sqrt{3}}{2};"

Answer: "z_1=\alpha;"

"z_2=-\dfrac{1}{2}\alpha+i\dfrac{\alpha\sqrt{3}}{2};"

"z_3=-\dfrac{1}{2}\alpha-i\dfrac{\alpha\sqrt{3}}{2};"